# Client-wise Modality Selection for Balanced Multi-modal Federated Learning

### v1

@@ 5.4 Ablation study
Effectiveness of each component. Tab. 1 studies the effect of each CMSFed component. Applying Eq. 9 (FedAvg+PCE) surpasses the vanilla strategy (FedAvg) by a large margin, demonstrating its effectiveness on local enhancement. Comparing CMSFed (64.5% on IID CREMA-D) with ‘DivFL+PCE’ (57.1% on the same setting) also denotes the necessity of balancing different modalities via modality selection. To show the importance of aligning feature spaces of weak modality, we replace the global prototypes with local prototypes (CMSFed-local). Global alignment achieves notable improvement (by up to 2% on non-IID AVE). The performance improvement compared with ‘FedAvg-0.8,0.5,0.2’ exhibits that randomly sampling modalities does not always lead to improvement and further demonstrates the need of meticulously selecting modalities for information exploitation.
@@
The relationship between modality selection and imbalance degree. In Eqs. 13 and 14, we use a conflict resolution strategy based on local imbalance ratio to realize balanced modality selection. We visualize the proportions of audio and visual modalities selected in each round and the global imbalance ratio. It is clear from Fig. 6 that audio is the dominant modality (imbalance ratio is always greater than 1) and modality imbalance is gradually alleviated as training progresses (global imbalance ratio shows a downward trend). In addition, the proportion of selected visual modality follows the same trend (larger in the early stage of training and becomes smaller later), implying the rationality of our selection strategy based on imbalance ratio and its effectiveness on mitigating bias.
Effectiveness on modality incongruity scenario. All above experiments assume that each client initially has complete modal data. Here, we build the stimulation of modality incongruity scenario, in which half of the clients have data with two modalities, and the other half only have data in one modality: random audio or visual. The results are shown in Tab. 2 (The results of pow-d is not available because it is not applicable to this scenario). Our CMSFed still makes impressive improvement compared with all other baselines (by up to 3.5% on IID CREMA-D), illustrating the good generalization ability of our method in different scenarios. It is worth mentioning that FedMSplit performs better than before because it is specifically designed for modality incongruity.
@@ 6 Conclusion
In this paper, we analyze traditional client selections in MFL and reveal there exists strong modality-level bias due to the modality imbalance during the training iterations. To address this issue, we propose the client-wise modality selection scheme for MFL (CMSFed) with modality-level gradient decoupling to release the potential of all modalities and maximize the gradient diversities to improve global aggregation. Our method does not introduce additional local training costs and communication overheads compared with previous methods. Extensive experiments on three datasets demonstrate the superiority of our method in performance and applicability under different modal combinations, data distributions and modality incongruity scenarios.

### v2

@@ 4.4 Ablation study
Effectiveness of each component. Tab. 4 studies the effect of each BMSFed component. Applying ME loss Eq. 7 on random sampling (FedAvg+ME) and the well-designed client selection (DivFL+ME) surpasses their vanilla strategies (FedAvg, DivFL) by a large margin, demonstrating its effectiveness on local enhancement. Comparing BMSFed (64.5% on IID CREMA-D) with ‘DivFL+ME’ (57.1% on the same setting) also denotes the necessity of balancing different modalities considering the global model via modality selection. To show the importance of aligning feature spaces of weak modality, we replace the global prototypes with local prototypes (BMSFed-local). Global alignment achieves notable improvement (by up to 2% on non-IID AVE). The performance improvement compared with ‘FedAvg-0.2,0.5,0.8’ exhibits that randomly sampling modalities does not always lead to improvement and further demonstrates the need of meticulously selecting modalities for information exploitation.
(a) CREMA-D under IID
(b) CREMA-D under IID
(c) CG-MNIST under IID
Figure 3: Robustness validation on data size, local epoch and client number. Our BMSFed consistently outperforms baseline (FedAvg) under various scenarios.
@@
Robustness test. To verify the robustness of our method, we vary three key hyperparameters to build various scenarios: (1) change the data size $\left|\mathcal{D}_{i}\right|$ to allow each client to hold a small amount of data, (2) set different local training epochs, and (3) vary the total client number $N$. As illustrated in Figs. 3(a) and 3(b), our BMSFed consistently outperforms baseline (FedAvg) under various scenarios. More data as well as more local training epochs can bring further improvements to our method, indicating that exploiting the weak modality need more training efforts. Based on the results in Fig. 3(c), our approach can also be generalised to clients with larger scales.
The relationship between modality selection and imbalance degree. In Eqs. 12 and 13, we use a conflict resolution strategy based on local imbalance ratio to realize balanced modality selection. We visualize the proportions of audio and visual modalities selected in each round and the global imbalance ratio. It is clear from Fig. 4 that audio is the dominant modality (imbalance ratio is always greater than 1) and modality imbalance is gradually alleviated as training progresses (global imbalance ratio shows a downward trend). In addition, the proportion of selected visual modality follows the same trend (larger in the early stage of training and becomes smaller later), implying the rationality of our selection strategy based on imbalance ratio and its effectiveness on mitigating bias.
Effectiveness on modality incongruity scenario. All above experiments assume that each client initially has complete modal data. Here, we build the stimulation of modality incongruity scenario, in which half of the clients have data with two modalities, and the other half only have data in one modality: random audio or visual. The results are shown in Tab. 5 (The results of pow-d is not available because it is not applicable to this scenario). Our BMSFed still makes impressive improvement compared with all other baselines (by up to 3.5% on IID CREMA-D), illustrating the good generalization ability of our method in different scenarios. It is worth mentioning that FedMSplit performs better than before because it is specifically designed for modality incongruity.
@@ 5 Conclusion
In this paper, we analyze traditional client selections and find their ineffectiveness in MFL. We further reveal that there exists strong modality-level bias due to the modality imbalance during the training iterations and uni-modal training on some clients may contribute more to the global model than multi-modal training. To address this issue, we propose the balanced modality selection scheme for MFL (BMSFed) with modality-level gradient decoupling to release the potential of all modalities and maximize the gradient diversities to improve global aggregation. We also introduce a modal enhancement loss to optimize the local update process. Our method does not introduce additional local training costs and communication overheads compared with previous methods. Extensive experiments on four datasets demonstrate the superiority of our method in performance and applicability under different modal combinations, data distributions and modality incongruity scenarios.
Acknowledgements
Our work was supported by two grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU15222621, PolyU15225023) and the funding from National Natural Science Foundation of China under grants 62302184.